# Shapley Values-enabled Progressive Pseudo Bag Augmentation for Whole-Slide Image Classification

### v2

@@ Abstract
In computational pathology, whole-slide image (WSI) classification presents a formidable challenge due to its gigapixel resolution and limited fine-grained annotations. Multiple-instance learning (MIL) offers a weakly supervised solution, yet refining instance-level information from bag-level labels remains complex. While most of the conventional MIL methods use attention scores to estimate instance importance scores (IIS) which contribute to the prediction of the slide labels, these often lead to skewed attention distributions and inaccuracies in identifying crucial instances. To address these issues, we propose a new approach inspired by cooperative game theory: employing Shapley values to assess each instance’s contribution, thereby improving IIS estimation. The computation of the Shapley value is then accelerated using attention, meanwhile retaining the enhanced instance identification and prioritization. We further introduce a framework for the progressive assignment of pseudo bags based on estimated IIS, encouraging more balanced attention distributions in MIL models. Our extensive experiments on CAMELYON-16, BRACS, and TCGA-LUNG datasets show our method’s superiority over existing state-of-the-art approaches, offering enhanced interpretability and class-wise insights. We will release the code upon acceptance.
@@ 1 Introduction
Recent advancements in digital pathology and artificial intelligence have significantly expanded the potential for analyzing whole-slide images (WSIs) in diagnostic contexts, prognostic evaluations, and various clinical tasks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. A key area within this domain is WSI classification [15, 16, 17, 18, 19, 20, 21, 22, 23], a fundamental and vital process distinguished by the gigapixel resolution of WSIs, setting it apart from typical natural image classification. The complex nature of the WSI classification task necessitates the adoption of specialized methodologies such as multi-instance learning (MIL) [24, 25, 26, 27, 28]. The principle of MIL is that the presence of at least one positive instance within a bag classifies the entire bag as positive; otherwise, it is considered negative. Most of the current research in MIL builds on the essential idea of distilling more instance-level information from bag-level labels [29, 30, 31, 32]. In this paradigm, attention-based pooling [33] stands out as a prominent technique, as the attention score $\alpha$ it generates for each instance in the bag naturally serves as a choice for estimating the contribution of each instance, referred to as the instance importance score (IIS). For example, attention scores play a crucial role in assisting MIL models in discerning significant instances to mitigate overfitting [34, 35, 36, 37]. Moreover, some studies leverage attention scores to fine-tune the feature encoder [38, 39, 40]. The attention score is so powerful that these studies all operate under the assumption that crucial (positive) instances can be identified by selecting those with top-ranking attention scores.
@@
To address these two inherent problems within attention-based MIL, in this study, we propose a progressive pseudo bag augmented MIL framework, termed PMIL. This framework takes full advantage of pseudo bag augmentation under the guidance of the Shapley value. Specifically, we first apply pseudo bag augmentation to MIL, aiming at encouraging models to focus on more important instances. Furthermore, to improve the mislabeling issue in pseudo bag augmentation, we introduce the Shapley value as a means of IIS estimation to constrain the assignment strategy instead of random splitting. Our approach divides a regular bag into a series of pseudo bags in a reasonable manner, thereby reducing the intrinsic noise associated with pseudo bag creation and enhancing the model’s generalization ability. In summary, our main contributions are as follows:
@@
  • Extensive experiments on the CAMELYON-16, BRACS, and TCGA-LUNG datasets demonstrate that our method outperformed other state-of-the-art methods in both slide-level and instance-level evaluation, and provided class-wise interpretation with Shapley value-based IIS.
@@ 2.1 Multiple Instance Learning for WSI Classification
2.2 Pseudo Bag Augmentation for MIL
Addressing the challenge of extreme attention distribution, a notable strategy involves dividing the regular bag into several pseudo bags [43, 44]. This approach encourages MIL models to learn from a greater variety of bags. Existing pseudo bag-based approaches adopted a random splitting strategy. For instance, DTFD [45] proposed a double-tier MIL model upon features distilled from pseudo bags to solve the mislabeling issue. As depicted in Fig. 1(c), DTFD smoothed out the attention distribution compared to other non-pseudo-bag-augmented MIL methods, which verifies the efficacy of pseudo bag augmentation to some extent. PseMix [46] applied a bag-prototype-based clustering to constrain the random pseudo bag assignment, and introduced the MixUp strategy for training. However, these methods fail to fundamentally solve the inherent mislabeling issue associated with random pseudo bag assignment.
2.3 Shapley Value Approximation
In the context of the misidentification of positive instances, the Shapley value offers a solution by quantifying the contribution of each instance based on its interactions with others [47, 48, 49]. Javed et al. [50] applied the Shapley value as an alternative to attention scores for MIL training and inference, considering it as the weighting score of each feature. However, this approach falls short of addressing the substantial computational demands of calculating Shapley values. Given their exponential complexity, these calculations often become prohibitively resource-intensive, posing a significant challenge in practical applications.
@@ 3.2 Shapley Value-based IIS Estimation
where $x_{i,j}$ is the $j$-th feature in the $i$-th bag to calculate Shapley value $\phi_{i,j}$, $X_{i}$ is the full feature set of the $i$-th bag, $S_{i}\subseteq X_{i}\backslash\left\{x_{i,j}\right\}$ are all available subsets.
@@
Table 1: Bag-level performance results for 5 repeatability experiments on CAMELYON-16, BRACS, and TCGA-LUNG test set. The subscripts are the standard deviation of each metric. The best evaluation results are in bold.
Method CAMELYON-16 BRACS TCGA-LUNG Average over Three Datasets
@@
MeanMIL 68.41.8 72.61.9 61.63.3 52.42.6 69.21.6 40.62.5 82.00.9 88.92.0 82.01.0 68.4 72.6 61.6
MaxMIL 76.11.8 85.72.6 74.45.1 55.92.8 75.91.6 50.34.0 88.71.0 94.41.2 88.71.0 76.1 85.7 74.4
ABMIL [33] 76.21.9 84.32.1 74.31.7 58.40.9 76.10.6 54.72.3 87.60.7 93.11.8 87.60.7 76.2 84.3 74.3
DSMIL [35] 72.21.7 77.12.1 68.92.6 53.12.2 70.83.3 46.13.7 86.21.4 93.61.0 86.21.4 72.2 77.1 68.9
CLAM [34] 74.83.2 81.62.4 73.33.5 53.83.5 73.31.7 51.53.3 88.21.4 94.21.2 88.21.4 74.8 81.6 73.3
TransMIL [42] 76.51.4 86.20.7 73.91.3 57.02.4 75.51.0 49.25.2 87.90.9 94.80.8 87.90.9 76.5 86.2 73.9
DTFD [45] 77.81.6 77.33.2 76.31.7 57.22.7 76.62.0 56.23.8 88.80.6 94.60.8 88.80.6 77.8 77.3 76.3
PMIL 87.41.1 90.11.6 86.31.1 67.13.3 82.81.8 66.43.1 91.31.4 96.50.9 91.31.4 81.3 89.8 81.3
@@ 4.1 Datasets and Evaluation Metrics
Our experimental setup employs three publicly available datasets to assess the performance of our proposed method.
@@ 4.2 Implementation Details
In the preprocessing stage, we utilize OTSU’s thresholding method to detect and localize tissue regions for patch generation. We create non-overlapping patches measuring 256×256 pixels at magnifications of 20× for CAMELYON-16 and TCGA-LUNG, and 5× for BRACS. This process results in an average of approximately 7156, 11951, and 714 patches per bag for these datasets, respectively.
All experiments were performed on a workstation equipped with NVIDIA RTX 3090 GPUs. We employed ResNet50 [57] as the encoder and ABMIL [33] as the primary MIL model. The Adam optimizer, with a weight decay of 1e-5, was selected. We also implemented an early stopping strategy, setting the patience parameter to 20 epochs. The initial learning rate was established at 3e-4 and subsequently reduced to 1e-4 for fine-tuning purposes. For the CAMELYON-16 dataset, we limited the maximum number of pseudo bags to 8; for BRACS, the limit was 10; and for TCGA-LUNG, it was 14. The increment in the number of pseudo bags is set to 4. In terms of Shapley value computation acceleration, we set the parameter $\mu$ to 10 and $\tau$ to 3. The total EM training round $n$ is set to 10.
@@ 4.3 Evaluation and Comparison
We present the experimental results of our proposed PMIL framework built on the ABMIL [33] backbone for CAMELYON-16, BRACS, and TCGA-LUNG datasets, comparing them with the following methods: (1) Conventional instance-level MIL, including the Mean-Pooling MIL and Max-Pooling MIL. (2) The vanilla attention-based MIL, ABMIL [33]. (3) Two variants of ABMIL, including non-local attention pooling DSMIL [35], single-attention-branch CLAM-SB [34]. (4) transformer-based MIL, TransMIL [42]. (5) Pseudo bag augmented MIL, DTFD [45].
Figure 4: Heatmaps of 4 slide sub-fields using different models and IIS estimations. (a) and (b) are macro and micro metastasis cases from CAMELYON-16, where red annotations are cancer regions and green annotations are noncancer regions in the column of ’Ground Truth’. (c) is the malignant tumor case from BRACS, where blue annotations are the malignant tumor (MT) regions and red annotations are atypical tumor (AT) regions in the column of ’Ground Truth’. In other columns, brighter red colors indicate higher probabilities.
Table 3: Performance results of pseudo bag augmentation using different IIS measure metrics on CAMELYON-16, BRACS, and TCGA-LUNG test sets. The subscripts are the standard deviation. The best evaluation results are in bold.
Metrics CAMELYON-16 BRACS TCGA-LUNG
ACC AUC F1 ACC AUC F1 ACC AUC F1
Random $87.13_{0.9}$ $86.19_{2.5}$ $85.49_{1.2}$ $61.78_{2.4}$ $80.39_{0.5}$ $58.46_{5.9}$ $89.65_{1.9}$ $95.77_{1.0}$ $89.64_{1.9}$
Attention Score $\textbf{87.44}_{2.8}$ $89.76_{0.7}$ $86.28_{2.7}$ $\textbf{68.28}_{1.7}$ $\textbf{83.98}_{0.3}$ $\textbf{66.47}_{2.4}$ $90.33_{1.3}$ $95.57_{0.8}$ $90.31_{1.3}$
Shapley Value $\textbf{87.44}_{1.1}$ $\textbf{90.10}_{1.6}$ $\textbf{86.30}_{1.1}$ $67.13_{3.3}$ $82.82_{1.8}$ $66.42_{3.1}$ $\textbf{91.29}_{1.2}$ $\textbf{96.45}_{0.8}$ $\textbf{91.29}_{1.2}$
In the bag-level evaluation, as shown in Table 1, our proposed PMIL framework demonstrates remarkable performance, achieving AUC scores of 90.1% for CAMELYON-16, 82.8% for BRACS, and 96.5% for TCGA-LUNG. These scores consistently exceed those of all other methods included in the comparison. Notably, on the complex BRACS dataset, our approach exhibits significant superiority. The generation of progressively refined pseudo bags contributes to enhanced training diversity and a reduction in the number of instances per bag. This strategy effectively improves the proficiency of the model in learning from positive instances.
@@ 4.4 Visualization and Interpretation
To emphasize the limitations of the attention score-based IIS, we conducted a comparative analysis, as illustrated in Fig. 4(a) and (b). In cases of macro metastasis, both ABMIL and our model show effective performance. However, in micro metastasis scenarios, the attention score-based IIS suggests that both ABMIL and our model erroneously focus on some noncancerous areas, which eludes logical interpretation. Conversely, using Shapley value-based IIS, our model precisely excludes noncancerous regions and accurately pinpoints cancerous areas.
Distinct from attention scores, the computation of Shapley values encompasses the entire MIL classifier, incorporating diverse category information, thereby enabling interpretations on a class-wise basis. As shown in Fig. 4(c), the IIS estimated by attention scores and class-MT Shapley values predominantly focus on malignant tumor regions. Meanwhile, the heatmaps generated using class-AT Shapley values predominantly emphasize atypical tumor regions, aligning with the slide-level labels. Although the heatmaps might not be entirely accurate for the BRACS dataset, this finding highlights the robust interpretability of Shapley value-based IIS in multiple classification tasks.
@@ 4.5.1 IIS Measure Estimation Metrics
In this ablation study, we evaluated both the attention score and the Shapley value as methods for estimating IIS for subsequent training, using random splitting as the baseline for pseudo bag augmentation. According to the results presented in Table 3, Shapley value-based IIS estimation demonstrates superior performance on the CAMELYON-16 and TCGA-LUNG datasets. Conversely, on the BRACS dataset, the attention score-based estimation yielded better results. This variation in effectiveness is likely due to the direct acquisition of attention scores via pooling operations. In contrast, calculating the Shapley value requires an additional fully connected layer, which may be less robust when the overall performance of the MIL classifier is not particularly high. Therefore, the Shapley value estimation is more advantageous with datasets that pose fewer learning challenges.
Figure 5: AUC results of different maximum pseudo bag numbers $M_{max}$ on CAMELYON-16, BRACS, and TCGA-LUNG datasets.
Table 4: AUC Results of the proposed method using different hyper-parameters on CAMELYON-16, and BRACS test sets.
AUC Dataset
Camelyon16 BRACS
$\mu$ $\tau$ Avg Std Avg Std
5 3 90.51 0.57 82.38 1.84
10 90.10 1.61 82.82 1.84
15 90.00 0.93 82.61 1.50
20 90.40 1.16 82.96 1.02
10 1 90.33 1.26 82.29 1.24
2 90.94 0.97 82.64 0.76
3 90.10 1.61 82.82 1.84
4 89.52 0.61 82.65 1.77
5 91.04 1.55 82.89 1.13
@@
Pseudo Bag Strategy CAMELYON-16 BRACS TCGA-LUNG
Number Initialization ACC AUC F1 ACC AUC F1 ACC AUC F1
Constant Constant 80.821.0 77.511.7 76.931.6 62.651.3 82.780.9 60.891.8 90.571.4 95.920.3 90.551.4
Progressive Constant 84.892.8 85.481.6 82.413.8 64.662.1 82.250.8 62.122.6 89.742.3 95.551.6 89.722.3
Constant Progressive 85.084.2 86.414.1 83.335.4 70.690.6 84.490.7 68.710.8 90.570.3 96.060.6 90.560.3
Progressive Progressive 88.181.2 88.101.0 86.991.5 71.261.2 84.880.2 69.861.4 91.301.7 96.091.3 91.281.8

### v4

@@ Abstract
In computational pathology, whole-slide image (WSI) classification presents a formidable challenge due to its gigapixel resolution and limited fine-grained annotations. Multiple-instance learning (MIL) offers a weakly supervised solution, yet refining instance-level information from bag-level labels remains challenging. While most of the conventional MIL methods use attention scores to estimate instance importance scores (IIS) which contribute to the prediction of the slide labels, these often lead to skewed attention distributions and inaccuracies in identifying crucial instances. To address these issues, we propose a new approach inspired by cooperative game theory: employing Shapley values to assess each instance’s contribution, thereby improving IIS estimation. The computation of the Shapley value is then accelerated using attention, meanwhile retaining the enhanced instance identification and prioritization. We further introduce a framework for the progressive assignment of pseudo bags based on estimated IIS, encouraging more balanced attention distributions in MIL models. Our extensive experiments on CAMELYON-16, BRACS, TCGA-LUNG, and TCGA-BRCA datasets show our method’s superiority over existing state-of-the-art approaches, offering enhanced interpretability and class-wise insights. Our source code is available at https://github.com/RenaoYan/PMIL.
@@ 1 Introduction
Recent advancements in digital pathology and artificial intelligence have significantly expanded the potential for analyzing whole-slide images (WSIs) in diagnostic contexts, prognostic evaluations, and various clinical tasks [1, 2, 3, 4, 5, 6, 7, 8]. A key area within this domain is WSI classification [9, 10, 11, 12, 13, 14], a fundamental and vital process distinguished by the gigapixel resolution of WSIs, setting it apart from typical natural image classification. The complex nature of the WSI classification task necessitates the adoption of specialized methodologies such as multi-instance learning (MIL) [15, 16, 17, 18, 19].
The principle of MIL is that the presence of at least one positive instance within a bag classifies the entire bag as positive; otherwise, it is considered negative. Most of the current research in MIL builds on the essential idea of distilling more instance-level information from bag-level labels [20, 21, 22, 23]. In this paradigm, attention-based pooling [24] stands out as a prominent technique, as the attention score $\alpha$ it generates for each instance in the bag naturally serves as a choice for estimating the contribution of each instance, referred to as the instance importance score (IIS). For example, attention scores play a crucial role in assisting MIL models in discerning significant instances to mitigate overfitting [25, 26, 27, 28]. Moreover, some studies leverage attention scores to fine-tune the feature encoder [29, 30, 31]. The attention score is so powerful that these studies all operate under the assumption that crucial (positive) instances can be identified by selecting those with top-ranking attention scores.
@@
To address these two inherent problems within attention-based MIL, in this study, we propose a progressive pseudo bag augmented MIL framework, termed PMIL. This framework takes full advantage of pseudo bag augmentation under the guidance of the Shapley value. Derived from the cooperative game [32], the Shapley value offers a solution by quantifying the contribution of each instance based on its interactions with others [33, 34]. Specifically, we first apply pseudo bag augmentation to MIL, aiming at encouraging models to focus on more important instances. Furthermore, to improve the mislabeling issue in pseudo bag augmentation, we introduce the Shapley value as a means of IIS estimation to constrain the assignment strategy instead of random splitting. Our approach divides a regular bag into a series of pseudo bags in a reasonable manner, thereby reducing the intrinsic noise associated with pseudo bag creation and enhancing the model’s generalization ability. In summary, our main contributions are as follows:
@@
  • Extensive experiments on the CAMELYON-16, BRACS, TCGA-LUNG, and TCGA-BRCA datasets demonstrate that our method outperformed other state-of-the-art methods in both slide-level and instance-level evaluation, and provided class-wise interpretation with Shapley values.
@@ 2.1 Multiple Instance Learning for WSI Classification
2.2 Advanced IIS related techniques for MIL
Addressing the challenge of extreme attention distribution, a notable strategy involves dividing the regular bag into several pseudo bags [37, 38]. This approach encourages MIL models to learn from a greater variety of bags. Existing pseudo bag-based approaches adopted a random splitting strategy. For instance, DTFD [39] proposed a double-tier MIL model upon features distilled from pseudo bags to solve the mislabeling issue. As depicted in Fig. 1(c), DTFD smoothed out the attention distribution compared to other non-pseudo-bag-augmented MIL methods, which verifies the efficacy of pseudo bag augmentation to some extent. PseMix [40] applied a bag-prototype-based clustering to constrain the random pseudo bag assignment, and then introduced the MixUp strategy to assign mixed pseudo bags with soft targets for training.
In the context of the misidentification of positive instances, Javed et al. [41] applied the Shapley value as an alternative to attention scores for MIL training and inference, considering it as the weighting score of each feature.
In summary, current pseudo bag augmented MIL methods typically employ a random splitting strategy, which inevitably introduces mislabeling noise into both training and inference phases. Additionally, their pseudo bag assignment is static and cannot be learned during training. Although Javed et al. [41] were the first to introduce the Shapley value to MIL, they did not address the significant computational demands of calculating Shapley values, which presents a major challenge for practical applications. Furthermore, their code lacks detailed training procedures, making it difficult to reproduce their performance results.
@@ 3.2 Shapley Value-based IIS Estimation
where $x_{i,j}$ is the $j$-th feature in the $i$-th bag to calculate the Shapley value $\phi_{i,j}$, $X_{i}$ is the full feature set of the $i$-th bag, $S_{i}\subseteq X_{i}\backslash\left\{x_{i,j}\right\}$ are all available subsets.
@@
Table 1: Bag-level performance results for 5 repeatability experiments on CAMELYON-16, BRACS, TCGA-LUNG, and TCGA-BRCA test sets. The subscripts are the standard deviation of each metric. The best evaluation results are in bold.
Method CAMELYON-16 BRACS TCGA-LUNG TCGA-BRCA
@@
MeanMIL 70.91.8 58.71.9 62.33.3 52.42.6 69.21.6 40.62.5 82.00.9 88.92.0 82.01.0 84.02.2 84.75.1 72.15.7
MaxMIL 83.71.8 86.72.6 83.35.1 55.92.8 75.91.6 50.34.0 88.71.0 94.41.2 88.71.0 84.51.8 84.64.7 70.46.3
ABMIL [24] 82.51.9 83.82.1 80.61.7 58.40.9 76.10.6 54.72.3 87.60.7 93.11.8 87.60.7 84.92.3 82.74.8 75.83.6
DSMIL [26] 77.21.7 77.22.1 74.42.6 53.12.2 70.83.3 46.13.7 86.21.4 93.61.0 86.21.4 84.41.1 83.25.2 73.43.9
CLAM [25] 82.53.2 81.62.4 80.13.5 53.83.5 73.31.7 51.53.3 88.21.4 94.21.2 88.21.4 85.92.8 85.84.2 77.05.3
TransMIL [36] 85.01.4 89.10.7 83.31.3 57.02.4 75.51.0 49.25.2 87.90.9 94.80.8 87.90.9 83.42.1 83.44.3 68.16.5
DTFD [39] 85.31.6 85.43.2 84.91.7 57.22.7 76.62.0 56.23.8 88.80.6 94.60.8 88.80.6 85.62.3 84.64.1 76.44.1
PseMix [40] 85.63.3 86.14.4 78.35.5 57.93.5 77.31.9 53.25.6 89.80.6 95.21.3 89.50.9 84.91.6 84.43.6 61.52.7
Additive MIL [41] 65.36.1 56.43.4 59.64.8 54.32.6 74.20.5 47.41.9 80.93.6 88.31.4 80.83.8 83.51.9 81.84.4 72.44.0
PMIL 87.41.1 90.11.6 86.31.1 67.13.3 82.81.8 66.43.1 91.31.4 96.50.9 91.31.4 86.91.5 86.34.2 79.03.0
@@ 4.1 Datasets and Evaluation Metrics
Our experimental setup employs four publicly available datasets to assess the performance of our proposed method.
@@
TCGA-BRCA comprises 985 WSIs, encompassing 797 invasive ductal carcinoma (IDL) and 198 invasive lobular carcinoma (ILC) cases. We adopt a 5-fold cross-validation protocol for both training and testing.
@@ 4.2 Implementation Details
In the preprocessing stage, we utilize OTSU’s thresholding method to detect and localize tissue regions for patch generation. We create non-overlapping patches measuring 256$\times$256 pixels at magnifications of 20$\times$ for CAMELYON-16, TCGA-LUNG, and TCGA-BRCA, and 5$\times$ for BRACS. This process results in an average of approximately 7156, 11951, 2392, and 714 patches per bag for these datasets, respectively.
All experiments were performed on a workstation equipped with NVIDIA RTX 3090 GPUs. We employed ResNet50 [48] as the encoder and ABMIL [24] as the primary MIL model. The Adam optimizer, with a weight decay of 1e-5, was selected. We also implemented an early stopping strategy, setting the patience parameter to 20 epochs. The initial learning rate was established at 3e-4 and subsequently reduced to 1e-4 for fine-tuning purposes. For the CAMELYON-16 dataset, we limited the maximum number of pseudo bags to 8; for BRACS, TCGA-LUNG, and TCGA-BRCA, the limits were 10, 14, and 6, respectively. The increment in the number of pseudo bags is set to 4. In terms of Shapley value computation acceleration, we set the parameter $\mu$ to 10 and $\tau$ to 3. The total EM training round $n$ is set to 10.
@@ 4.3 Evaluation and Comparison
We present the experimental results of our proposed PMIL framework built on the ABMIL [24] backbone for CAMELYON-16, BRACS, TCGA-LUNG, and TCGA-BRCA datasets, comparing them with the following methods: (1) Conventional instance-level MIL, including the Mean-Pooling MIL and Max-Pooling MIL. (2) The vanilla attention-based MIL, ABMIL [24]. (3) Two variants of ABMIL, including non-local attention pooling, DSMIL [26], and single-attention-branch, CLAM-SB [25]. (4) Transformer-based MIL, TransMIL [36]. (5) Pseudo bag augmented MIL, DTFD [39] and PseMix [40]. (6) Shapley value enabled MIL, Additive MIL.
Figure 4: Heatmaps of two slide sub-fields using different models. (a) and (b) are macro and micro metastasis cases from CAMELYON-16, where red and green annotations are cancer and noncancer regions in the column of ’Ground Truth’.
Figure 5: Heatmaps of one malignant tumor (MT) case from BRACS with various IIS estimations. The blue and red annotations are the malignant tumor and atypical tumor (AT) regions in the ’Ground Truth’.
In the bag-level evaluation, as shown in Table 1, our proposed PMIL framework demonstrates remarkable performance, achieving AUC scores of 90.1% for CAMELYON-16, 82.8% for BRACS, 96.5% for TCGA-LUNG, and 86.3% for TCGA-BRCA. These scores consistently exceed those of all other methods included in the comparison. Notably, on the complex BRACS dataset, our approach exhibits significant superiority. The generation of progressively refined pseudo bags contributes to enhanced training diversity and a reduction in the number of instances per bag. This strategy effectively improves the proficiency of the model in learning from positive instances.
@@ 4.4 Visualization and Interpretation
To emphasize the limitations of the attention score-based IIS, we conducted a comparative analysis, as illustrated in Fig. 4(a) and (b). In cases of macro metastasis, ABMIL, DTFD, and our model show effective performance. However, in micro metastasis scenarios, the attention score-based IIS suggests that ABMIL, DTFD, and our model erroneously focus on some noncancerous areas, which eludes logical interpretation. Conversely, using Shapley value-based IIS, our model precisely excludes noncancerous regions and accurately pinpoints cancerous areas.
Figure 6: Ablation results of different maximum pseudo bag numbers $M_{max}$ on CAMELYON-16, BRACS, TCGA-LUNG, and TCGA-BRCA datasets.
Table 3: Performance results of pseudo bag augmented MIL using different IIS measurements on CAMELYON-16, BRACS, TCGA-LUNG, and TCGA-BRCA test sets. The subscripts are the standard deviation. The best evaluation results are in bold.
Metrics CAMELYON-16 BRACS TCGA-LUNG TCGA-BRCA
ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
[…]
Distinct from attention scores, the computation of Shapley values encompasses the entire MIL classifier, incorporating diverse category information, thereby enabling interpretations on a class-wise basis. As shown in Fig. 5, the IIS estimated by attention scores and class-MT Shapley values predominantly focuses on malignant tumor regions. Meanwhile, the heatmaps generated using class-AT Shapley values predominantly emphasize atypical tumor regions, aligning with the slide-level labels. Although the heatmaps might not be entirely accurate for the BRACS dataset, this finding highlights the robust interpretability of Shapley value-based IIS in multiple classification tasks.
@@ 4.5.1 IIS Measure Estimation Metrics
In this ablation study, we evaluated both the attention score and the Shapley value as methods for estimating IIS for subsequent training, using random splitting as the baseline for pseudo bag augmentation. According to the results presented in Table 3, Shapley value-based IIS estimation demonstrates superior performance on the CAMELYON-16, TCGA-LUNG, and TCGA-BRCA datasets. Conversely, on the BRACS dataset, the attention score-based estimation yielded better results. This variation in effectiveness is likely due to the direct acquisition of attention scores via pooling operations. In contrast, calculating the Shapley value requires an additional fully connected layer, which may be less robust when the overall performance of the MIL classifier is not particularly high. Therefore, the Shapley value estimation is more advantageous with datasets that pose fewer learning challenges.
Table 4: AUC results of the proposed method using different hyper-parameters on CAMELYON-16, BRACS, TCGA-LUNG and TCGA-BRCA test sets. The subscripts are the standard deviation. The best evaluation results are in bold.
$\mu$ $\tau$ CAMELYON-16 BRACS TCGA-LUNG TCGA-BRCA
5 3 90.510.57 82.381.84 95.480.14 85.314.99
10 90.101.61 82.821.84 95.740.27 86.914.35
15 90.000.93 82.611.50 95.710.53 86.494.64
20 90.401.16 82.961.02 95.690.48 86.383.30
10 1 90.331.26 82.291.24 95.690.71 85.353.40
2 90.940.97 82.640.76 95.570.95 87.074.78
3 90.101.61 82.821.84 95.740.27 86.914.35
4 89.520.61 82.651.77 96.260.88 85.923.78
5 91.041.55 82.891.13 96.080.32 85.676.43
@@
Pseudo Bag Strategy CAMELYON-16 BRACS TCGA-LUNG TCGA-BRCA
Number Initialization ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
Constant Constant 80.821.0 77.511.7 76.931.6 62.651.3 82.780.9 60.891.8 90.571.4 95.920.3 90.551.4 85.482.0 85.803.4 73.906.0
Progressive Constant 84.892.8 85.481.6 82.413.8 64.662.1 82.250.8 62.122.6 89.742.3 95.551.6 89.722.3 86.501.7 85.893.7 77.883.6
Constant Progressive 85.084.2 86.414.1 83.335.4 70.690.6 84.490.7 68.710.8 90.570.3 96.060.6 90.560.3 86.292.1 85.833.9 77.673.3
Progressive Progressive 88.181.2 88.101.0 86.991.5 71.261.2 84.880.2 69.861.4 91.301.7 96.091.3 91.281.8 86.801.7 86.034.3 78.653.0